OpenStudy (anonymous):

solve for the variable indicated 9w=4x for w

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

@JayBaybie

OpenStudy (lyrae):

If you are to solve for w, your goal is to transfer the 9 from the left side to the right side. Do you kow how to do that?

OpenStudy (anonymous):

no

OpenStudy (lyrae):

Whats \[\frac{ 9 }{ 9 } \]?

OpenStudy (anonymous):

1

OpenStudy (lyrae):

So if you divide the left side by 9 you'll have \[\frac{ 9 }{ 9 }w = 1 w = w\] What operation do you have to perform on the right side to keep both sides balanced?

OpenStudy (anonymous):

i don't know

OpenStudy (lyrae):

The same we did on the left side! Which means we also have to devide the right side by 9. What does that become?

OpenStudy (anonymous):

9

OpenStudy (lyrae):

It's \[\frac{ 4x }{ 9 }\] So this means we have \[9w = 4x \rightarrow \frac{ 9w }{ 9 } = \frac{ 4x }{ 9 } \rightarrow w = \frac{ 4x }{ 9 }\]Where the last equation is the solution to your problem.

OpenStudy (anonymous):

thank you